OpenStudy (anonymous):

Someone Please help Which pair of ratios form a proportion? A. 3 : 4 and 4 : 5 B. 10 : 12 and 16 : 18 C. 7 : 10 and 10 : 14 D. 5 : 8 and 20 : 32

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (mathteacher1729):

This is asking "Which pairs of ratios reduce to the same fraction?" Yes or no; 3/4 = 5/6 ? 10/12 = 16/18? 7/10 = 10/14? 5/8 = 20/32? Reduce all fractions, only one of those choices will result in a true equality.